OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5

OpenStudy (anonymous):

you want to first distribute everything

OpenStudy (anonymous):

then get all the "x" values to one side and all the numbers to the other

OpenStudy (anonymous):

−3(x + 3) = −3(x + 1) − 5 x+3=(x+1)+5/3 divide both sides by -3

OpenStudy (anonymous):

looks simpler now.

OpenStudy (anonymous):

ugh can you just like do it cuz im really bad at this.. :/

OpenStudy (anonymous):

−3(x + 3) = −3(x + 1) − 5 x+3=(x+1)+5/3 divide both sides by -3 x=x+1+(5/3)-3 subtract 3 from both sides x=x-1/3 add like terms the 1+(5/3)-3 are like terms It is false

OpenStudy (anonymous):

as you can see by final result x=x-1/3 pick any number lets say 1 for x 1=1-1/3 it is false 2=2-1/3 lets say x=2 then it is also false.